Find the equation of each line. Write each answer in slope-intercept form.
Example
The line with slope -8 that goes through (-1, -5)

Answer by matthew_sessoms(39)

Since we know the POINT and the SLOPE...obviously, lets use the POINT-SLOPE formula (or form).


y - y1 = m(x - x1)


y + 5 = -8(x + 1)
y = -8x - 8 - 5
y = -8x - 13


-Matthew